# Physics of droplet regulation in biological cells

### v3

@@ 3.1 Complex molecular interactions
Interactions between biomolecules can be roughly divided into weak, transient interactions and strong, persistent bonds. While weak multivalent interactions are generally responsible for phase separation into liquid-like phases, strong interactions tend to form long-lived assemblies, which can persist as nano-domains inside condensates [237]. Moreover, chemical modifications of biomolecules can affect their interactions. For instance, proteins can be changed by post-translational modifications [199], which are implicated in condensate regulation [238, 54, 239, 36]. Such changes can induce multimerisation, which can in turn enable phase separation. To understand this phenomenon, we imagine a molecule whose weak interactions are insufficient for phase separation. A multimer of such molecules, however, can interact with many other multimers, which effectively increases the molecular volume $\Omega$ while maintaining the interaction strength $\chi$ per unit volume [172]. This process essentially reduces the entropic cost of concentrating molecules, promoting phase separation. This feature has been used to induce artificial condensate in cells [14, 240]. More generally, it is likely that combinations of weak and strong interactions enable fine-tuning of condensate properties [241, 242].
@@ 3.2.1 Internal stresses
For a single-component Newtonian fluid, the total stress can be decomposed into contributions arising from hydrostatic pressure and viscous dissipation [121]. The hydrostatic pressure $p$ enforces incompressibility of the fluid, and is determined by momentum conservation (Navier–Stokes; equation 2.9) and the incompressibility condition $\nabla\cdot\boldsymbol{v}=0$. Multi-component fluids additionally induce an ‘equilibrium stress’, $\boldsymbol{\sigma}_{\textrm{eq}}$, that captures osmotic and interfacial effects. Further, many biological fluids exhibit a complex rheology that results in an elastic contribution to the stress, denoted by $\boldsymbol{\sigma}_{\textrm{el}}$. The overall stress tensor $\boldsymbol{\sigma}$ then comprises the sum of each of these different contributions,
@@ 3.4 Electrostatic effects
where the first term is given by equation (2.6), while the second term captures electrostatic effects. The associated electrochemical potential, $\mu_{\mathrm{e}}=\nu\delta F_{\mathrm{e}}/\delta\phi_{i}$, gains a term proportional to $\varphi$, $\mu_{\mathrm{e}}=\mu+q_{i}\varphi$, relative to the uncharged case. This electrochemical potential shows that positively charged molecules move down gradients in the electrostatic potential $\varphi$. In particular, small mobile ions will tend to distribute so that the system becomes charge neutral, inhibiting long-ranged electrostatic effects. In the simplest case, these ions can be described by Debye-Hückel theory, which accounts for their translational entropies and electrostatic interactions [314, 315]. The main result is that electrostatic effects do not extend significantly beyond the Debye length $\lambda_{\mathrm{D}}=[\varepsilon k_{\mathrm{B}}T/(\sum_{i}c_{i}q_{i}^{2})]^{1/2}$, where $c_{i}$ is the mean number concentration of the ion species $i$ with charge density $q_{i}$. Note that ion concentrations can be phase-dependent [296], meaning that electrostatic interactions may be longer ranged than would be expected from the average salt concentration [306].
@@
where the non-dimensional factor $\alpha=\frac{1}{3}(4\pi)^{2}\ell_{\mathrm{B}}^{3/2}\nu_{0}^{-1/2}$ governs the strength of electrostatic interactions. Here, the Bjerrum length $\ell_{\mathrm{B}}=e^{2}/(4\pi\varepsilon k_{\mathrm{B}}T)$ describes the distance beyond which thermal energy dominates electrostatic repulsion of two elementary charges $e$ [324]. A typical value is $\ell_{\mathrm{B}}\approx 0.7\,\mathrm{nm}$ since the permittivity $\varepsilon$ of cytosol is comparable to water [325]. We thus have $\alpha\sim 0.1\ldots 1$ for typical molecular volumes of $\nu_{0}\sim 10^{3}\ldots 10^{5}\,\mathrm{nm^{3}}$, and the electrostatic contribution becomes comparable to the enthalpic contributions in $f$ if $|q_{i}|\sim e$, indicating that electrostatic effects can lower the free energy significantly. The theory given by equation (3.6) qualitatively explains complex coacervation, but it cannot predict realistic coacervates quantitatively [326], particularly when the oppositely charged molecules are asymmetric [327, 304, 162, 305, 306, 328].
@@ 3.5 Multiphase coexistence
Figure 3.1: Multi-component phase diagrams. Phase diagrams of ${N_{\mathrm{c}}}=2$ solutes and an inert solvent as a function of the solute fractions $\phi_{1}$ and $\phi_{2}$ for two scenarios with different interaction matrices. A system with an average composition within the blue region can split into two coexisting phases whose compositions follow from the dashed tie lines.
@@ 4.1 Theory of wetting
where $f_{\mathrm{surf}}$ is the free energy density of surface interactions, also known as a contact potential. Minimising this extended free energy results in unchanged bulk chemical potentials (see appendix B), in addition to boundary conditions
@@ 4.1.1 Partial, full and pre- wetting
Figure 4.1: Different phenomena resulting from droplet–structure interactions. (A) Dense phase partially wets the external boundary, forming a compact droplet whose interface meets the boundary at a contact angle $\vartheta$. The areas of the interfaces between the droplet and the dilute phase, and the droplet and the boundary are denoted by $A$ and $S$, respectively. (B) Dense phase completely wets the boundary. (C) In an undersaturated system, droplet material pre-wets the boundary, forming a layer of thickness $h$. (D) Droplet volume fraction $\phi$ as a function of distance from the boundary. Note that the gradient of $\phi$ at $z=0$ is the same for all three cases. (E) Phase diagram showing the pre-, full and partial wetting regimes. The typical size of the pre-wetting regime has been exaggerated for clarity. (F) Droplet–droplet interactions between two immiscible phases result in either droplet separation, engulfment, or partial wetting (clockwise from top left). Panels (A) to (E) inspired by figures in [364] and panel (F) inspired by figure in [14].
@@ 4.1.4 Heterogeneous nucleation
This expression is obtained analogously to the energy barrier for homogeneous nucleation (equation 2.34) by finding the maximum value of $\Delta F$ prescribed by equation (4.4). For perfectly non-wetting conditions ($\vartheta=\pi$) the classic homogeneous nucleation result is recovered. Any other contact angle results in preferential nucleation at the surface. Analogous expressions can also obtained for more complex surface shapes [372]. By tuning the strength of interactions between droplet material and other cellular structures, cells are thus able to regulate the location and rate at which condensates are nucleated [373, 374, 375, 376].
@@ 5.1 Chemical reactions of many components
$k_{\alpha}(\Phi)=\hat{k}_{\alpha}\prod_{i=1}^{N_{\mathrm{c}}}\phi_{i}^{\frac{1}{2}\bigl{(}\sigma^{(\alpha)}_{\rightarrow,i}+\sigma^{(\alpha)}_{\leftarrow,i}\bigr{)}}\;$, (5.5)
@@
Chemical reactions also take place in heterogeneous systems, where the composition $\Phi$ varies in space and time. The general dynamics of such systems is discussed in section 2, where we already introduced the source term $s_{i}$ in equation (2.14) to account for chemical reactions. The combination of equation (2.14) with equation (5.4) for $s^{(\alpha)}$ implies that chemical reactions affect composition only locally, whereas spatial transport takes place via diffusion and advection.
@@ 5.2.1 Stability of homogeneous states
The influence of chemical reactions on the stability of the homogeneous system is more subtle (Figure 5.2). For autocatalytic reactions where $s^{\prime}(\phi_{0})>0$, the homogeneous state is always unstable since increasing $\phi$ further accelerates its production, destabilizing the overall composition, $\omega(\boldsymbol{q}=0)>0$. In contrast, the overall composition is stable $\omega(\boldsymbol{q}=0)<0$ for auto-inhibitory reactions where $s^{\prime}(\phi_{0})<0$. In this case, the fastest growing mode corresponds to a length scale $\ell_{\mathrm{max}}=2\pi/q_{\mathrm{max}}=\pi[-8\kappa/f^{\prime\prime}(\phi_{0})]^{1/2}$ and grows at rate $\omega_{\mathrm{max}}=s^{\prime}(\phi_{0})+\Lambda[f^{\prime\prime}(\phi_{0})]^{2}/4\kappa$. Consequently, auto-inhibitory reactions only exhibit instabilities in the spinodal region where $f^{\prime\prime}(\phi_{0})<0$. However, even in the spinodal region, the homogeneous state can be completely stabilised if the reactions are sufficiently strong, $s^{\prime}(\phi_{0})<k_{\mathrm{stab}}$, where $k_{\mathrm{stab}}=-\Lambda[f^{\prime\prime}(\phi_{0})]^{2}/4\kappa$. For weaker reactions, $k_{\mathrm{stab}}<s^{\prime}(\phi_{0})<0$, only a compact band of wavelengths is unstable (Figure 5.2). If $f^{\prime\prime}(\phi_{0})<0$, the length scale $\ell_{\mathrm{max}}$ of the most unstable mode is independent of $s^{\prime}(\phi_{0})$, so we expect standard spinodal decomposition to initially develop in the intermediate regime. In particular, droplets will form and Ostwald ripening (see section 2.2.4) may drive coarsening. However, the fact that $\omega$ is negative for long wavelengths (small $\boldsymbol{q}$) in this intermediate regime suggests that late stage dynamics may deviate from the classical ripening behaviour. To investigate this possibility, we next analyse the dynamics of multiple droplets.
@@ 5.2.2 Effective droplet model
where $\phi_{\infty}=\phi^{(0)}_{\mathrm{out}}+\Gamma_{\mathrm{out}}/k_{\mathrm{out}}$ is the volume fraction far away from the droplet and the equilibrium fraction $\phi^{\mathrm{eq}}_{\mathrm{out}}$ is given by equation (2.32). Equation (5.10) differs in two crucial aspects from equation (2.39) describing passive droplets. First, the fraction $\phi_{\infty}$ is now a constant value controlled by the reactions in the dilute phase, in contrast to the passive case where it was determined by material conservation and depended on the size of all droplets. Indeed the dynamics of multiple droplets are independently described by equation (5.10), as long as the droplet separation is large enough compared to $\ell_{\mathrm{out}}$ such that they do not disturb their corresponding concentration profiles. Second, equation (5.10) contains a term related to the reactions inside the droplet, which can dominate growth over the surface tension effects captured by the other terms in this equation.
@@ 5.3.3 Suppressed nucleation
to linear order in $k$. The pre-factor here is identical to the passive energy barrier given by equation (2.34). Reactions thus increase $\Delta\!F_{\mathrm{nucl}}$, implying a smaller nucleation rate $J\propto\exp(-\Delta\!F_{\mathrm{nucl}})$; see equation (2.35). Since chemical reactions can be interpreted as a separation of charge, this suppression of nucleation follows from the fact that forming a charged nucleus is energetically unfavourable compared to forming a neutral nucleus. This process can also be understood from the fact that chemical reactions want to maintain a fixed concentration locally and thus dampen fluctuations originating from diffusive fluxes, making the formation of a critical nucleus less likely.
Chemical reactions also affect heterogeneous droplet nucleation at surfaces [465]. In section 4.1.4, we discuss how surfaces promote nucleation, and this behaviour is retained when reactions describing externally maintained droplets are added [465]. However, sufficiently large reaction rates will still suppress nucleation, as for the homogeneous case. Further, reactions may cause deformation of sessile droplets at surfaces (see figure 5.4b), as discussed in more detail in section 5.5.1 below. The interactions of droplets with both surfaces and active reactions thus provide flexible control over droplet nucleation.
@@ 5.4 Internally maintained droplets
Internally maintained droplets of radius $R$ exhibit an efflux of droplet material $J\approx 4\pi DR\left(\phi^{\mathrm{eq}}_{\mathrm{in}}-\phi_{\infty}\right)$ (see figure 5.3B), whose functional form follows from solving equation (5.9) in the stationary state. Here, $\phi^{\mathrm{eq}}_{\mathrm{in}}=\phi^{(0)}_{\mathrm{in}}(1+\ell_{\gamma}/R)$ is the equilibrium fraction inside the droplet and $\phi_{\infty}=\phi^{(0)}_{\mathrm{out}}+\Gamma_{\mathrm{out}}/k_{\mathrm{out}}=\phi^{(0)}_{\mathrm{out}}(1+\mathcal{S})$ denotes the volume fraction in the dilute phase. To nucleate a droplet, the efflux $J$ needs to be compensated for by local production of droplet material, which could originate from localised enzymes, e.g., at the surface of a catalytically active core of radius $a$. Since the integrated reaction flux $Q$ at to such a core needs to exceed $J$, there is a minimal flux $Q_{\mathrm{min}}$ beyond which an internally maintained droplet is nucleated. For weak surface tension effects ($\ell_{\gamma}\ll a$), we find $Q_{\mathrm{min}}\approx 4\pi Da(\phi^{(0)}_{\mathrm{in}}-\phi_{\infty})$, whereas strong surface tension ($\ell_{\gamma}\gg a$) implies $Q_{\mathrm{min}}\approx 4\pi D\ell_{\gamma}\phi^{(0)}_{\mathrm{out}}$. In both cases, nucleation of internally maintained droplets around the core is ensured, while spontaneous nucleation anywhere else in the system is virtually impossible.

### v4

@@ 3.1 Complex molecular interactions
Interactions between biomolecules can be roughly divided into weak, transient interactions and strong, persistent bonds. While weak multivalent interactions are generally responsible for phase separation into liquid-like phases, strong interactions tend to form long-lived assemblies, which can persist as nano-domains inside condensates [237]. Moreover, chemical modifications of biomolecules can affect their interactions. For instance, proteins can be changed by post-translational modifications [199], which are implicated in condensate regulation [238, 54, 239, 36]. Such changes can induce multimerisation, which can in turn enable phase separation. To understand this phenomenon, we imagine a molecule whose weak interactions are insufficient for phase separation. A multimer of such molecules, however, can interact with many other multimers, which effectively increases the molecular volume $\Omega$ while maintaining the interaction strength $\chi$ per unit volume [172]. This process essentially reduces the entropic cost of concentrating molecules, promoting phase separation. This feature has been used to induce artificial condensates in cells [14, 240]. More generally, it is likely that combinations of weak and strong interactions enable fine-tuning of condensate properties [241, 242].
@@ 3.2.1 Internal stresses
For a single-component Newtonian fluid, the total stress can be decomposed into contributions arising from hydrostatic pressure and viscous dissipation [121]. The hydrostatic pressure $p$ enforces incompressibility of the fluid, and is determined by momentum conservation (Navier–Stokes; equation 2.9) and the incompressibility condition $\nabla\cdot\boldsymbol{v}=0$. Multi-component fluids additionally induce an equilibrium stress, $\boldsymbol{\sigma}_{\textrm{eq}}$, that captures osmotic and interfacial effects. Further, many biological fluids exhibit a complex rheology that results in an elastic contribution to the stress, denoted by $\boldsymbol{\sigma}_{\textrm{el}}$. The overall stress tensor $\boldsymbol{\sigma}$ then comprises the sum of each of these different contributions,
@@ 3.4 Electrostatic effects
where the first term is given by equation (2.6), while the second term captures electrostatic effects. The associated electrochemical potential, $\mu_{i}^{\mathrm{e}}=\nu\delta F_{\mathrm{e}}/\delta\phi_{i}$, gains a term proportional to $\varphi$, $\mu_{i}^{\mathrm{e}}=\mu_{i}+q_{i}\varphi$, relative to the uncharged case. This electrochemical potential shows that positively charged molecules move down gradients in the electrostatic potential $\varphi$. In particular, small mobile ions will tend to distribute so that the system becomes charge neutral, inhibiting long-ranged electrostatic effects. In the simplest case, these ions can be described by Debye-Hückel theory, which accounts for their translational entropies and electrostatic interactions [314, 315]. The main result is that electrostatic effects do not extend significantly beyond the Debye length $\lambda_{\mathrm{D}}=[\varepsilon k_{\mathrm{B}}T/(\sum_{i}c_{i}q_{i}^{2})]^{1/2}$, where $c_{i}$ is the mean number concentration of the ion species $i$ with charge density $q_{i}$. Note that ion concentrations can be phase-dependent [296], meaning that electrostatic interactions may be longer ranged than would be expected from the average salt concentration [306].
@@
where the non-dimensional factor $\alpha=\frac{1}{3}(4\pi)^{2}\ell_{\mathrm{B}}^{3/2}\nu_{0}^{-1/2}$ governs the strength of electrostatic interactions. Here, the Bjerrum length $\ell_{\mathrm{B}}=e^{2}/(4\pi\varepsilon k_{\mathrm{B}}T)$ describes the distance beyond which thermal energy dominates electrostatic repulsion of two elementary charges $e$ [324]. A typical value is $\ell_{\mathrm{B}}\approx 0.7\,\mathrm{nm}$ since the permittivity $\varepsilon$ of cytosol is comparable to water [325]. We thus have $\alpha\sim 1\ldots 100$ for typical molecular volumes of $\nu_{0}\sim 10^{-1}\ldots 10^{3}\,\mathrm{nm^{3}}$, and the electrostatic contribution becomes larger than enthalpic contributions in $f$ if $|q_{i}|\sim e$, indicating that electrostatic effects can lower the free energy significantly. The theory given by equation (3.6) qualitatively explains complex coacervation, but it cannot predict realistic coacervates quantitatively [326], particularly when the oppositely charged molecules are asymmetric [327, 304, 162, 305, 306, 328].
@@ 3.5 Multiphase coexistence
Figure 3.1: Multi-component phase diagrams. Phase diagrams of ${N_{\mathrm{c}}}=2$ solutes and an inert solvent as a function of the solute fractions $\phi_{1}$ and $\phi_{2}$ for two scenarios with different interaction matrices $\chi_{ij}$. A system with an average composition within the blue region can split into two coexisting phases whose compositions follow from the dashed tie lines.
@@ 4.1 Theory of wetting
where $f_{\mathrm{surf}}$ is the free energy density of surface interactions, also known as a contact potential. Minimising this extended free energy results in unchanged bulk chemical potentials (appendix B), in addition to boundary conditions
@@ 4.1.1 Partial, full and pre- wetting
Figure 4.1: Different phenomena resulting from droplet–structure interactions. (A) Dense phase partially wets the external boundary, forming a compact droplet whose interface meets the boundary at a contact angle $\vartheta$. The areas of the interfaces between the droplet and the dilute phase, and the droplet and the boundary are denoted by $A$ and $S$, respectively. (B) Dense phase completely wets the boundary. (C) In an undersaturated system, droplet material pre-wets the boundary, forming a layer of thickness $h$. (D) Droplet volume fraction $\phi$ as a function of distance $z$ from the boundary. Note that the gradient of $\phi$ at $z=0$ is the same for all three cases. (E) Phase diagram showing the pre-, full and partial wetting regimes. The typical size of the pre-wetting regime has been exaggerated for clarity. (F) Droplet–droplet interactions between two immiscible phases result in either droplet separation, engulfment, or partial wetting (clockwise from top left). Panels (A) to (E) inspired by figures in [364] and panel (F) inspired by figure in [14].
@@ 4.1.4 Heterogeneous nucleation
This expression is obtained analogously to the energy barrier for homogeneous nucleation (equation 2.34) by finding the maximum value of $\Delta F$ prescribed by equation (4.4). For perfectly non-wetting conditions ($\vartheta=\pi$) the classic homogeneous nucleation result is recovered. Any other contact angle results in preferential nucleation at the surface. Analogous expressions can also be obtained for more complex surface shapes [372]. By tuning the strength of interactions between droplet material and other cellular structures, cells are thus able to regulate the location and rate at which condensates are nucleated [373, 374, 375, 376].
@@ 5.1 Chemical reactions of many components
$k_{\alpha}(\Phi)=\hat{k}_{\alpha}\prod_{i=1}^{N_{\mathrm{c}}}\phi_{i}^{\frac{1}{2}\bigl(\sigma^{(\alpha)}_{\rightarrow,i}+\sigma^{(\alpha)}_{\leftarrow,i}\bigr)}$ (5.5)
@@
Chemical reactions also take place in heterogeneous systems, where the composition $\Phi$ varies in space and time. The general dynamics of such systems are discussed in section 2, where we already introduced the source term $s_{i}$ in equation (2.14) to account for chemical reactions. The combination of equation (2.14) with equation (5.4) for $s^{(\alpha)}$ implies that chemical reactions affect composition only locally, whereas spatial transport takes place via diffusion and advection.
@@ 5.2.1 Stability of homogeneous states
The influence of chemical reactions on the stability of the homogeneous system is more subtle (Figure 5.2). For autocatalytic reactions where $s^{\prime}(\phi_{0})>0$, the homogeneous state is always unstable since increasing $\phi$ further accelerates its production, destabilizing the overall composition, $\omega(\boldsymbol{q}=0)>0$. In contrast, the overall composition is stable $\omega(\boldsymbol{q}=0)<0$ for auto-inhibitory reactions where $s^{\prime}(\phi_{0})<0$. In this case, the fastest growing mode corresponds to a length scale $\ell_{\mathrm{max}}=2\pi/q_{\mathrm{max}}=\pi[-8\kappa/f^{\prime\prime}(\phi_{0})]^{1/2}$ and grows at rate $\omega_{\mathrm{max}}=s^{\prime}(\phi_{0})+\Lambda[f^{\prime\prime}(\phi_{0})]^{2}/4\kappa$. Consequently, auto-inhibitory reactions only exhibit instabilities in the spinodal region where $f^{\prime\prime}(\phi_{0})<0$. However, even in the spinodal region, the homogeneous state can be completely stabilised if the reactions are sufficiently strong, $s^{\prime}(\phi_{0})<k_{\mathrm{stab}}$, where $k_{\mathrm{stab}}=-\Lambda[f^{\prime\prime}(\phi_{0})]^{2}/4\kappa$. For weaker reactions, $k_{\mathrm{stab}}<s^{\prime}(\phi_{0})<0$, only a compact band of wavelengths is unstable (Figure 5.2). If $f^{\prime\prime}(\phi_{0})<0$, the length scale $\ell_{\mathrm{max}}$ of the most unstable mode is independent of $s^{\prime}(\phi_{0})$, so we expect standard spinodal decomposition to initially develop in the intermediate regime. In particular, droplets will form and Ostwald ripening (section 2.2.4) may drive coarsening. However, the fact that $\omega$ is negative for long wavelengths (small $\boldsymbol{q}$) in this intermediate regime suggests that late stage dynamics may deviate from the classical ripening behaviour. To investigate this possibility, we next analyse the dynamics of multiple droplets.
@@ 5.2.2 Effective droplet model
where $\phi_{\infty}=\phi^{(0)}_{\mathrm{out}}+\Gamma_{\mathrm{out}}/k_{\mathrm{out}}$ is the volume fraction far away from the droplet and the equilibrium fraction $\phi^{\mathrm{eq}}_{\mathrm{out}}$ is given by equation (2.32). Equation (5.10) differs in two crucial aspects from equation (2.39) describing passive droplets. First, the fraction $\phi_{\infty}$ is now a constant value controlled by the reactions in the dilute phase, in contrast to the passive case where it was determined by material conservation and depended on the size of all droplets. Indeed, the dynamics of multiple droplets are independently described by equation (5.10), as long as the droplet separation is large enough compared to $\ell_{\mathrm{out}}$ such that they do not disturb their corresponding concentration profiles. Second, equation (5.10) contains a term related to the reactions inside the droplet, which can dominate growth over the surface tension effects captured by the other terms in this equation.
@@ 5.3.3 Suppressed nucleation
to linear order in $k$. The pre-factor here is identical to the passive energy barrier given by equation (2.34). Reactions thus increase $\Delta\!F_{\mathrm{nucl}}$, implying a smaller nucleation rate $J_{\mathrm{nucl}}\propto\exp(-\Delta\!F_{\mathrm{nucl}}/k_{\mathrm{B}}T)$; see equation (2.35). Since chemical reactions can be interpreted as a separation of charge, this suppression of nucleation follows from the fact that forming a charged nucleus is energetically unfavourable compared to forming a neutral nucleus. This process can also be understood from the fact that chemical reactions want to maintain a fixed concentration locally and thus dampen fluctuations originating from diffusive fluxes, making the formation of a critical nucleus less likely.
Chemical reactions also affect heterogeneous droplet nucleation at surfaces [465]. In section 4.1.4, we discuss how surfaces promote nucleation, and this behaviour is retained when reactions describing externally maintained droplets are added [465]. However, sufficiently large reaction rates will still suppress nucleation, as for the homogeneous case. Further, reactions may cause deformation of sessile droplets at surfaces (figure 5.4B), as discussed in more detail in section 5.5.1 below. The interactions of droplets with both surfaces and active reactions thus provide flexible control over droplet nucleation.
@@ 5.4 Internally maintained droplets
Internally maintained droplets of radius $R$ exhibit an efflux of droplet material $J\approx 4\pi DR\left(\phi^{\mathrm{eq}}_{\mathrm{in}}-\phi_{\infty}\right)$ (figure 5.3B), whose functional form follows from solving equation (5.9) in the stationary state. Here, $\phi^{\mathrm{eq}}_{\mathrm{in}}=\phi^{(0)}_{\mathrm{in}}(1+\ell_{\gamma}/R)$ is the equilibrium fraction inside the droplet and $\phi_{\infty}=\phi^{(0)}_{\mathrm{out}}+\Gamma_{\mathrm{out}}/k_{\mathrm{out}}=\phi^{(0)}_{\mathrm{out}}(1+\mathcal{S})$ denotes the volume fraction in the dilute phase. To nucleate a droplet, the efflux $J$ needs to be compensated for by local production of droplet material, which could originate from localised enzymes, e.g., at the surface of a catalytically active core of radius $a$. Since the integrated reaction flux $Q$ at such a core needs to exceed $J$, there is a minimal flux $Q_{\mathrm{min}}$ beyond which an internally maintained droplet is nucleated. For weak surface tension effects ($\ell_{\gamma}\ll a$), we find $Q_{\mathrm{min}}\approx 4\pi Da(\phi^{(0)}_{\mathrm{in}}-\phi_{\infty})$, whereas strong surface tension ($\ell_{\gamma}\gg a$) implies $Q_{\mathrm{min}}\approx 4\pi D\ell_{\gamma}\phi^{(0)}_{\mathrm{out}}$. In both cases, nucleation of internally maintained droplets around the core is ensured, while spontaneous nucleation anywhere else in the system is virtually impossible.